# Wall in the stability space of the gluing stability conditions on Hirzebruch surfaces

### v1

@@ 3.2. Gluing perversity
${\rm per}(\sigma)=\phi(\lambda_{1}(\mathcal{O}_{x}))-\phi(\rho_{2}(\mathcal{O}_{x}))$ (3.18)
@@
$\phi(\lambda_{1}(\mathcal{O}_{x}))=\left\{\begin{aligned} &1&{\rm if\ \sigma_{1}\ is\ standard}\\ &\psi(Z_{1}(\mathcal{O}(k))+Z_{1}(\mathcal{O}(k-1)[1])&{\rm if\ \sigma_{1}\ is\ algebraic}\end{aligned}\right$. (3.19)
@@
$\phi(\rho_{2}(\mathcal{O}_{x}))=\left\{\begin{aligned} &1&{\rm if\ \sigma_{2}\ is\ standard}\\ &\psi(Z_{2}(\mathcal{O}(k))+Z_{2}(\mathcal{O}(k-1)[1])&{\rm if\ \sigma_{2}\ is\ algebraic}\end{aligned}\right$. (3.20)
@@
$p^{*}\mathcal{O}(k)[j]\to\mathcal{O}_{f}[j]\to p^{*}\mathcal{O}(k-1)[j+1].$ (3.21)
@@
${\rm per}_{i}(\sigma):=\left\{\begin{aligned} &1&{\rm if\ }\sigma_{i}{\rm\ is\ standard}\\ &\phi(Z_{i}(\mathcal{O}(k)))-\phi(Z_{i}(\mathcal{O}(k)[j+1]))&\quad{\rm if\ }\sigma_{i}{\rm\ is\ quiver}\end{aligned}\right$. (3.22)
@@
$\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}=\begin{pmatrix}-k&1&k+\frac{e}{2}&-1\\ 1-k&1&k-1+\frac{e}{2}&-1\\ 0&0&-k-\frac{e}{2}&1\\ 0&0&1-k-\frac{e}{2}&1\end{pmatrix}\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}$, (3.23)
$\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}=\begin{pmatrix}-1&1&-1&1\\ 1-k&k&1-k&k\\ 0&0&-1&1\\ 0&0&1-k-\frac{e}{2}&k+\frac{e}{2}\end{pmatrix}\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}$. (3.24)
@@
$\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}=\begin{pmatrix}0&0&k+\frac{e}{2}&-1\\ 0&0&k-1+\frac{e}{2}&-1\\ -k-e&1&-k-\frac{e}{2}&1\\ 1-k-e&1&1-k-\frac{e}{2}&1\end{pmatrix}\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}$, (3.25)
$\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}=\begin{pmatrix}-1&1&-1&1\\ 1-k-e&k+e&1-k-e&k+e\\ 1&-1&0&0\\ k-1+\frac{e}{2}&-k-\frac{e}{2}&0&0\end{pmatrix}\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}$. (3.26)
@@
$0\to\mathcal{O}_{f}\to\mathcal{O}_{x}\to\mathcal{O}_{f}(-C_{0})[1]\to 0$ (3.28)
@@
In general, for a smooth projective surface $S$ over $\mathbb{C}$, if $\mathcal{U}$ is the set of geometric stability condition of $\textrm{Stab}_{\Lambda}(S)$ (i.e. the skyscraper sheaves are stable of the same phase), then it is a well-known fact that $\mathcal{U}$ is an open subset of $\textrm{Stab}_{\Lambda}(S)$ ([Bri08]). Consequentry, the boundary $\partial\mathcal{U}:=\overline{\mathcal{U}}\setminus\mathcal{U}$ coincides with the set of stability conditions where all skyscraper sheaves are strictly $\sigma$-semistable. From 3.19, for gluing stability conditions, this strictly $\sigma$-semistability is equivalent to the condition ${\rm per}(\sigma)=0$.
@@
$\textrm{Re}Z_{gl}(\lambda_{1}(\mathcal{O}_{x}))\textrm{Im}Z_{gl}(\rho_{2}(\mathcal{O}_{x}))-\textrm{Im}Z_{gl}(\lambda_{1}(\mathcal{O}_{x}))\textrm{Re}Z_{gl}(\rho_{2}(\mathcal{O}_{x}))=0.$ (3.30)
@@ 4. Relation to divisorial stability conditions
$B\in\textrm{NS}(\Sigma_{e})_{\mathbb{R}},\quad\omega\in\textrm{Amp}(\Sigma_{e})_{\mathbb{R}}.$ (4.1)
@@
$\sigma_{\omega,B}=(Z_{\omega,B},\mathcal{A}_{\omega,B})$ (4.2)
@@
$\textrm{Amp}(S)_{\mathbb{R}}\times\textrm{NS}(S)_{\mathbb{R}}\hookrightarrow\textrm{Stab}_{\Lambda}(S)$ (4.3)
@@
$\mathcal{\pi}:\textrm{Stab}_{\Lambda}(S)\to\textrm{Hom}_{\mathbb{Z}}(\Lambda,\mathbb{C})$ (4.4)
@@
$\langle E,E^{\prime}\rangle_{M}=c_{1}.c_{1}^{\prime}-r^{\prime}\textrm{ch}_{2}-r\textrm{ch}_{2}^{\prime}.$ (4.5)
@@
$Z(E)=\langle\pi(\sigma),\textrm{ch}(E)\rangle_{M}.$ (4.6)
@@
$\partial\overline{S_{div}}=\partial_{z}\cup\partial_{w}.$ (4.8)
@@
$Z_{gl,1}=(-1)(-d_{\lambda_{1}}(E)+ir_{\lambda_{1}}(E))+(d_{\rho_{2}}-kr_{\rho_{2}})\zeta_{0}+(d_{\rho_{2}}+(1-k)r_{\rho_{2}})\zeta_{1}$ (4.9)
$Z_{gl,2}=-d_{\rho_{2}}(E)+ir_{\rho_{2}}(E)+(-1)((d_{\lambda_{1}}-kr_{\lambda_{1}})\zeta_{0}+(d_{\lambda_{1}}+(1-k)r_{\lambda_{1}})\zeta_{1})$ (4.10)
$Z_{gl,3}=(-1)(d_{\lambda_{1}}-kr_{\lambda_{1}})\zeta_{0}$ $+(d_{\lambda_{1}}+(1-k)r_{\lambda_{1}})\zeta_{1})$ (4.11)
$+(d_{\rho_{2}}-kr_{\rho_{2}})\zeta^{\prime}_{0}+(d_{\rho_{2}}+(1-k)r_{\rho_{2}})\zeta^{\prime}_{1}.$ (4.12)
@@
$Z_{gl,m}=\langle\pi(\sigma),\textrm{ch}\rangle_{M}.$ (4.13)
@@
$\pi(\sigma)M=\left(\begin{pmatrix}a\textrm{Re}\xi_{i}+b\textrm{Im}\xi_{i}\\ c\textrm{Re}\xi_{i}+d\textrm{Im}\xi_{i}\end{pmatrix}\right)_{i=0,1,2,3}.$ (4.14)
@@
$e^{B+i\omega}=\left(\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}x\\ z\end{pmatrix}C_{0},\begin{pmatrix}y\\ w\end{pmatrix}f,\begin{pmatrix}\frac{1}{2}((z^{2}-x^{2})e+2(xy-zw))\\ yz+xw-xze\end{pmatrix}\textrm{pt}\right).$ (4.15)
@@
$z=t\cdot\textrm{det}\begin{pmatrix}\textrm{Re}\xi_{0}&\textrm{Re}\xi_{1}\\ \textrm{Im}\xi_{0}&\textrm{Im}\xi_{1}\end{pmatrix}$, (4.16)
$w=t\cdot\textrm{det}\begin{pmatrix}\textrm{Re}\xi_{0}&\textrm{Re}\xi_{2}\\ \textrm{Im}\xi_{0}&\textrm{Im}\xi_{2}\end{pmatrix}$. (4.17)
@@ 4.3. The case of $m=3$
Since $c>0$ (as established earlier), we know $c+1\neq 0$. This implies, $\textrm{det}(\zeta_{0},\zeta_{1})-\textrm{det}(\zeta_{0}+\zeta_{1},\zeta^{\prime}_{0})=0$, which is the desired formula.
@@ 5. Moduli spaces of $\sigma_{gl}$-stable objects
${\rm gr}_{0}(\mathcal{O}_{x})=\mathcal{O}_{\Sigma_{e}}(kf)\oplus\mathcal{O}_{\Sigma_{e}}((k-1)f)[1]\oplus\mathcal{O}_{\Sigma_{e}}(-C_{0}+kf)[1]\oplus\mathcal{O}_{\Sigma_{e}}(-C_{0}+(k-1)f)[2]$ (5.2)
@@
$\textrm{ch}_{0}({\rm gr}_{0}(\mathcal{O}_{x}))=1-1-1+1=0,$ (5.3)
$\textrm{ch}_{1}({\rm gr}_{0}(\mathcal{O}_{x}))=kf-(k-1)f-(-C_{0}+kf)+(-C_{0}+(k-1)f)=0,$ (5.4)
$\textrm{ch}_{2}({\rm gr}_{0}(\mathcal{O}_{x}))=\frac{2k+e}{2}+\frac{-2k+2-e}{2}=1$ (5.5)

### v2

@@ 3.3. Gluing perversity
(3.28) ${\rm per}(\sigma)=\phi(\lambda_{1}(\mathcal{O}_{x}))-\phi(\rho_{2}(\mathcal{O}_{x}))$
@@
(3.29) $\phi(\lambda_{1}(\mathcal{O}_{x}))=\left\{\begin{aligned} &1&{\rm if\ \sigma_{1}\ is\ standard}\\ &\psi(Z_{1}(\mathcal{O}(k))+Z_{1}(\mathcal{O}(k-1)[1])&{\rm if\ \sigma_{1}\ is\ algebraic}\end{aligned}\right$.
@@
(3.30) $\phi(\rho_{2}(\mathcal{O}_{x}))=\left\{\begin{aligned} &1&{\rm if\ \sigma_{2}\ is\ standard}\\ &\psi(Z_{2}(\mathcal{O}(k))+Z_{2}(\mathcal{O}(k-1)[1])&{\rm if\ \sigma_{2}\ is\ algebraic}\end{aligned}\right$.
@@
(3.31) $p^{*}\mathcal{O}(k)[j]\to\mathcal{O}_{f}[j]\to p^{*}\mathcal{O}(k-1)[j+1].$
@@
(3.32) ${\rm per}_{i}(\sigma):=\left\{\begin{aligned} &1&{\rm if\ }\sigma_{i}{\rm\ is\ standard}\\ &\phi(Z_{i}(\mathcal{O}(k)))-\phi(Z_{i}(\mathcal{O}(k)[j+1]))&\quad{\rm if\ }\sigma_{i}{\rm\ is\ quiver}\end{aligned}\right$.
@@
(3.33) $\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}=\begin{pmatrix}-k&1&k+\frac{e}{2}&-1\\ 1-k&1&k-1+\frac{e}{2}&-1\\ 0&0&-k-\frac{e}{2}&1\\ 0&0&1-k-\frac{e}{2}&1\end{pmatrix}\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}$,
(3.34) $\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}=\begin{pmatrix}-1&1&-1&1\\ 1-k&k&1-k&k\\ 0&0&-1&1\\ 0&0&1-k-\frac{e}{2}&k+\frac{e}{2}\end{pmatrix}\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}$.
@@
(3.35) $\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}=\begin{pmatrix}0&0&k+\frac{e}{2}&-1\\ 0&0&k-1+\frac{e}{2}&-1\\ -k-e&1&-k-\frac{e}{2}&1\\ 1-k-e&1&1-k-\frac{e}{2}&1\end{pmatrix}\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}$,
(3.36) $\begin{pmatrix}\textrm{rank}(v)\\ \textrm{deg}_{C_{0}}(v)\\ \textrm{deg}_{f}(v)\\ \textrm{ch}_{2}(v)\end{pmatrix}=\begin{pmatrix}-1&1&-1&1\\ 1-k-e&k+e&1-k-e&k+e\\ 1&-1&0&0\\ k-1+\frac{e}{2}&-k-\frac{e}{2}&0&0\end{pmatrix}\begin{pmatrix}m_{0}\\ m_{1}\\ m_{2}\\ m_{3}\end{pmatrix}$.
@@
(3.38) $0\to\mathcal{O}_{f}\to\mathcal{O}_{x}\to\mathcal{O}_{f}(-C_{0})[1]\to 0$
@@
In general, for a smooth projective surface $S$ over $\mathbb{C}$, if $\mathcal{U}$ is the set of geometric stability condition of $\textrm{Stab}_{\Lambda}(S)$ (i.e. the skyscraper sheaves are stable of the same phase), then it is a well-known fact that $\mathcal{U}$ is an open subset of $\textrm{Stab}_{\Lambda}(S)$ ([Bri08]). Consequentry, the boundary $\partial\mathcal{U}:=\overline{\mathcal{U}}\setminus\mathcal{U}$ coincides with the set of stability conditions where all skyscraper sheaves are strictly $\sigma$-semistable. From 3.22, for gluing stability conditions, this strictly $\sigma$-semistability is equivalent to the condition ${\rm per}(\sigma)=0$.
@@
(3.40) $\textrm{Re}Z_{gl}(\lambda_{1}(\mathcal{O}_{x}))\textrm{Im}Z_{gl}(\rho_{2}(\mathcal{O}_{x}))-\textrm{Im}Z_{gl}(\lambda_{1}(\mathcal{O}_{x}))\textrm{Re}Z_{gl}(\rho_{2}(\mathcal{O}_{x}))=0.$
@@ 4. Relation to divisorial stability conditions
(4.1) $B\in\textrm{NS}(\Sigma_{e})_{\mathbb{R}},\quad\omega\in\textrm{Amp}(\Sigma_{e})_{\mathbb{R}}.$
@@
(4.2) $\sigma_{\omega,B}=(Z_{\omega,B},\mathcal{A}_{\omega,B})$
@@
(4.3) $\textrm{Amp}(S)_{\mathbb{R}}\times\textrm{NS}(S)_{\mathbb{R}}\hookrightarrow\textrm{Stab}_{\Lambda}(S)$
@@
(4.4) $\mathcal{\pi}:\textrm{Stab}_{\Lambda}(S)\to\textrm{Hom}_{\mathbb{Z}}(\Lambda,\mathbb{C})$
@@
(4.5) $\langle E,E^{\prime}\rangle_{M}=c_{1}.c_{1}^{\prime}-r^{\prime}\textrm{ch}_{2}-r\textrm{ch}_{2}^{\prime}.$
@@
(4.6) $Z(E)=\langle\pi(\sigma),\textrm{ch}(E)\rangle_{M}.$
@@
(4.8) $\partial\overline{S_{div}}=\partial_{z}\cup\partial_{w}.$
@@
(4.9) $Z_{gl,1}=(-1)(-d_{\lambda_{1}}(E)+ir_{\lambda_{1}}(E))+(d_{\rho_{2}}-kr_{\rho_{2}})\zeta_{0}+(d_{\rho_{2}}+(1-k)r_{\rho_{2}})\zeta_{1}$
(4.10) $Z_{gl,2}=-d_{\rho_{2}}(E)+ir_{\rho_{2}}(E)+(-1)((d_{\lambda_{1}}-kr_{\lambda_{1}})\zeta_{0}+(d_{\lambda_{1}}+(1-k)r_{\lambda_{1}})\zeta_{1})$
(4.11) $Z_{gl,3}=(-1)(d_{\lambda_{1}}-kr_{\lambda_{1}})\zeta_{0}$ $+(d_{\lambda_{1}}+(1-k)r_{\lambda_{1}})\zeta_{1})$
(4.12) $+(d_{\rho_{2}}-kr_{\rho_{2}})\zeta^{\prime}_{0}+(d_{\rho_{2}}+(1-k)r_{\rho_{2}})\zeta^{\prime}_{1}.$
@@
(4.13) $Z_{gl,m}=\langle\pi(\sigma),\textrm{ch}\rangle_{M}.$
@@
(4.14) $\pi(\sigma)M=\left(\begin{pmatrix}a\textrm{Re}\xi_{i}+b\textrm{Im}\xi_{i}\\ c\textrm{Re}\xi_{i}+d\textrm{Im}\xi_{i}\end{pmatrix}\right)_{i=0,1,2,3}.$
@@
(4.15) $e^{B+i\omega}=\left(\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}x\\ z\end{pmatrix}C_{0},\begin{pmatrix}y\\ w\end{pmatrix}f,\begin{pmatrix}\frac{1}{2}((z^{2}-x^{2})e+2(xy-zw))\\ yz+xw-xze\end{pmatrix}\textrm{pt}\right).$
@@
(4.16) $z=t\cdot\textrm{det}\begin{pmatrix}\textrm{Re}\xi_{0}&\textrm{Re}\xi_{1}\\ \textrm{Im}\xi_{0}&\textrm{Im}\xi_{1}\end{pmatrix}$,
(4.17) $w=t\cdot\textrm{det}\begin{pmatrix}\textrm{Re}\xi_{0}&\textrm{Re}\xi_{2}\\ \textrm{Im}\xi_{0}&\textrm{Im}\xi_{2}\end{pmatrix}$.
@@ 4.3. The case of $m=3$
Since $c>0$ (as established earlier), we know $c+1\neq 0$. This implies, $\textrm{det}(\zeta_{0},\zeta_{1})-\textrm{det}(\zeta_{0}+\zeta_{1},\zeta^{\prime}_{0})=0$, which is the desired formula. ∎
@@ 5. Moduli spaces of $\sigma_{gl}$-stable objects
(5.2) ${\rm gr}_{0}(\mathcal{O}_{x})=\mathcal{O}_{\Sigma_{e}}(kf)\oplus\mathcal{O}_{\Sigma_{e}}((k-1)f)[1]\oplus\mathcal{O}_{\Sigma_{e}}(-C_{0}+kf)[1]\oplus\mathcal{O}_{\Sigma_{e}}(-C_{0}+(k-1)f)[2]$
@@
(5.3) $\textrm{ch}_{0}({\rm gr}_{0}(\mathcal{O}_{x}))=1-1-1+1=0,$
(5.4) $\textrm{ch}_{1}({\rm gr}_{0}(\mathcal{O}_{x}))=kf-(k-1)f-(-C_{0}+kf)+(-C_{0}+(k-1)f)=0,$
(5.5) $\textrm{ch}_{2}({\rm gr}_{0}(\mathcal{O}_{x}))=\frac{2k+e}{2}+\frac{-2k+2-e}{2}=1$